OpenStudy (anonymous):

I have a question but I am stuck on the factoring 1/2*b*(h+7)=22

OpenStudy (anonymous):

Define 'factoring' - because that isn't something I'd look to factorise.

OpenStudy (anonymous):

1/2*b*(h+7)=22 1/2(b^2+h7)=22 2*1/2(b^2+h7)=2*22 b^2+h7=44 b^2+h7-44=44-44 b^2+h7-44=0 (b- )(h+ )= this is where I am lost

OpenStudy (anonymous):

u cant solve for b and h in same equation

OpenStudy (anonymous):

Here I will give you the question. The base of a triangle is 7cm greater than the height. The area is 22 cm^2. I have to find the height and base of the triangle.

OpenStudy (anonymous):

Define b = h+7 (or h = b-7) , and go from there.

OpenStudy (anonymous):

i.e. you should be solving: \[\frac{1}{2} \cdot h \cdot (h+7) = 22 \text{ or } \frac{1}{2} \cdot b \cdot (b-7) \]

OpenStudy (anonymous):

The last one =22 too :@

OpenStudy (radar):

Post the quadratic equation you get when you follow through with INewton's suggestion.

OpenStudy (radar):

My suggestion is to use the first one:\[(1/2) \times h \times(h + 7)=22 cm sq\]

OpenStudy (anonymous):

1/2*h(h+7)=22 1/2(h^2+h7)=22 2*1/2(h^2+h7)=2*22 h^2+h7=44 h^2-h7=44=44-44 h^2+h7-44=0 h=9 b=35

OpenStudy (radar):

\[h ^{2}=7h-44=0\] is correct. Which method did you use to get the values of h? quadratic formula, or factor method?

OpenStudy (radar):

\[h ^{2}+7h-44=0\] I meant to type sri

OpenStudy (radar):

which method did you use to get 9 for h??

OpenStudy (anonymous):

i think factor

OpenStudy (radar):

9 doesn't go into 44 so something went wrong. Factor is the way to go. That equation will factor beautifully. Do you need some guidance in factoring that equation?

OpenStudy (anonymous):

yes please

OpenStudy (radar):

Well we are in luck in that the coefficient of h^2 is 1. Also we need to find to factors of 44 which will when combined equal 7. We have a hint since 44 is negative, one of these factors will positive and the other negative. Do you follow so far?

OpenStudy (radar):

It will start looking like this: (h- ?) (h +?) the ?*?=-44 and ?+?=-7

OpenStudy (radar):

Any questions??? so far.

OpenStudy (radar):

leahpl81872 can you think of two numbers when multiplied equal -44 and when added =+7 ? It is important cause one of these numbers is your answer!

OpenStudy (anonymous):

11 & 4

OpenStudy (radar):

Right on!! (h+11)(h-4)=0

OpenStudy (radar):

which one do you think is the answer? the -11 or the 4 ???

OpenStudy (anonymous):

4

OpenStudy (radar):

very good, now remember the base was equal to the height + 7 cm, so what is the base?

OpenStudy (anonymous):

11

OpenStudy (radar):

Correcto, you seem to have a handle on this problem, good luck.

OpenStudy (anonymous):

thank you very much

OpenStudy (radar):

No problem, and INewton's input was helpful.